OpenStudy (anonymous):

(x-2/x+2)=(3/x) solv equation by any method, check each solution

OpenStudy (anonymous):

x(x-2) = 3(x+2) x^2 - 2x = 3x + 6 x^2 - 5x -6 = 0 x^2 - 6x + x -6=0 x(x-6) + 1(x-6)=0 (x+1)(x-6)=0 x= -1 or 6